# Super-extensions of tensor algebras and their applications

### v1

@@ Abstract.
Following [It12] and [It15], we construct two super-extensions of the usual tensor algebra through the super-actions of symmetric groups and Hecke algebras respectively. For each extension, we consider a special type of derivations coming from covectors, and study the the space generated, in some special manner, by these derivations and operators from left multiplication by vectors and permutations. Duality theorems of these spaces and the super-actions are proved. As an application, we provide a new proof of the Schur-Sergeev duality theorem, as well as its quantum version.
@@ 1. Introduction
Denote by $V^{\otimes d}$ the $d$-fold of a vector space $V$. In [Sc27], Schur considered the relation between the permutation action of the symmetric group $\mathfrak{S}_{d}$ and the diagonal action of the general linear group $\mathrm{GL(V)}$ on $V^{\otimes d}$. This is the well-known Schur-Weyl duality nowadays. It plays a fundamental role in representation theory (see, e.g., [We46]). Since its importance, this duality theorem has been generalized in several directions during the past decades. For example, in the late 1980s, Sergeev [Ser85] and Berele and Regev [BR87] discovered independently a super version of the Schur-Weyl duality. Meanwhile, Jimbo [Jim86] introduced a duality theorem about Hecke algebras and quantum enveloping algebras. In [Mi06], Mitsuhashi extended Jimbo’s theorem to the super case.
@@
Usually, in order to prove these dualities, people use the double commutant theorem. For example, for the classical Schur-Weyl duality, one needs to show that $\mathrm{GL}_{d}(\mathbb{C})=\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$. Then by the semisimplicity of the group algebra $\mathbb{C}\mathfrak{S}_{d}$ one has $\mathbb{C}\mathfrak{S}_{d}=\mathrm{End}_{\mathrm{GL}_{d}(\mathbb{C})}(\mathbb{C}^{d})$. The verification of the inclusion $\mathrm{GL}_{d}(\mathbb{C})\subset\mathrm{End}_{\mathfrak{S}_{d}}(\mathbb{C}^{d})$ is trivial. But the reverse direction requires further non-trivial techniques. We refer the reader to [GW03] and [CW12] for detailed proof.
@@
In [It12], Itoh generalized the usual tensor algebra $T(V)$ by considering actions of symmetric groups. He studied operators of left multiplication by vectors and linear functionals, and established a duality theorem between the space generated by these operators and the infinity symmetric group. As a consequence, he provided a new proof of the Schur-Weyl dual theorem. In a subsequent paper [It15], he generalized his constructions to quantum case in which symmetric group actions are replaced by Hecke algebra actions. Using a similar argument, he gave a new proof of Jimbo’s duality theorem.
The main purpose of this paper is to generalize Itoh’s results to super cases. First of all, we consider the right super-action of $\mathfrak{S}_{d}$ on $d$-fold of a vector superspace $V=V_{\overline{0}}\oplus V_{\overline{1}}$. Using this action, we define an associative algebra structure on the space
$\overline{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathfrak{S}_{\infty}.$
For each covector $v^{\ast}\in V^{\ast}$, we define a derivation $L(v^{\ast})$ of $\overline{T}(V)$. Denote by $L(w)$ and $L(\sigma)$ the operators coming from left multiplication by a vector $w$ and a permutation $\sigma$ respectively. We show that the space $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ spanned by operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast}),$
is the commutant of $\mathfrak{S}_{d+k}$. Notice that $\mathcal{A}_{d}^{0}$ is just action of the Lie superalgebra $\mathfrak{gl}(m|n)$. We provide a new proof of the Schur-Sergeev duality. Secondly, we consider the right super-action of the Hecke algebra $\mathscr{H}_{d}$ on $V^{\otimes d}$, and construct an algebra
$\widetilde{T}(V)=\bigoplus_{d\geq 0}V^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty}.$
Similar duality in this algebra is proved. As an application, we provide a new proof of Mitsuhashi’s duality.
@@
This paper is organized as follows. We construct the super-extension $\overline{T}(V)$ in Section 2 and study the communication relations among $L(v^{\ast})$, $L(w)$, and $L(\sigma)$ in Section 3. Duality theorems are proved in Section 4. We deal with the the quantum version $\widetilde{T}(V)$ in Section 5, 6, and 7.
@@ Acknowledgement
We would like to thank Li Luo for many helpful discussions. Especially, he introduced the Schur-Sergeev duality and its quantum version to us which initiate this work.
@@ 2.1. Symmetric groups
For any positive integer $d$, we denote by $\mathfrak{S}_{d}$ the symmetric group of the set $[d]=\{1,2,\ldots,d\}$, and by $s_{i}$, $11\leq i\leq d-$ 1, the transposition interchanging $i$ and $i+1$. We often use cycle notation to denote a permutation. So $\tau=(i_{1}\ i_{2}\ \cdots\ i_{k})\in\mathfrak{S}_{d}$ maps $i_{r}$ to $i_{r+1}$ for $11\leq r\leq k-$ 1 and $i_{r}$ to $i_{1}$, and fixes other numbers. There is a natural sequence of inclusions: $\mathfrak{S}_{1}\subset\cdots\subset\mathfrak{S}_{d}\subset\cdots$. The direct limit is denoted by $\mathfrak{S}_{\infty}$. One can view $\mathfrak{S}_{\infty}$ as the group generated by $\{s_{i}\}_{i\geq 1}$ subject to the relations:
@@
Given $k\geq 1$, we define a group homomorphism :${}^{\uparrow k}:\mathfrak{S}_{\infty}\rightarrow\mathfrak{S}_{\infty}$ by $s_{i}^{\uparrow k}=s_{i+k}$.
@@ 2.2. Superspaces and general linear Lie superalgebras
Let $m,n$ be two fixed positive integers, and $V=\mathbb{C}^{m|n}=\mathbb{C}^{m}\oplus\mathbb{C}^{n}$. As usual, we set $V_{\overline{0}}=\mathbb{C}^{m}$ and $V_{\overline{1}}=\mathbb{C}^{n}$. Vectors in $V_{\overline{0}}$ (resp. $V_{\overline{1}}$) are of parity 00 (resp. parity $1$). The parity of a homogeneous element $v\in V$ is denoted by $|v|$. In this section, we will denote by $e_{i}$, $1\leq i\leq m+n$, the vector in $V$ that is 1 in the $i$-th position and 0 elsewhere. Then $\{e_{1},\ldots,e_{m}\}$ and $\{e_{m+1},\ldots,e_{m+n}\}$ are bases of $V_{\overline{0}}$ and $V_{\overline{1}}$ respectively. We denote by $\{e_{1}^{\ast},\ldots,e_{m+n}^{\ast}\}$ the dual basis of $\{e_{1},\ldots,e_{m+n}\}$, i.e., $e_{i}^{\ast}(e_{j})=\delta_{ij}$. Here $\delta_{ij}$ is the Kronecker symbol. Define a parity function on this dual basis by requiring that $|e_{i}^{\ast}|=|e_{i}|$ for each $1\leq i\leq m+n$. Thus the dual space $V^{\ast}=V_{\overline{0}}^{\ast}\oplus V_{\overline{1}}^{\ast}$ of $V$ is a $\mathbb{Z}_{2}$-graded vector space.
@@
$\bigg{\{}\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}}.$
@@
$\mathfrak{gl}(m|n)_{\overline{0}}=\bigg{\{}\left(\begin{array}[]{cc}A&0\\ 0&D\end{array}\right)\bigg{|}A\in M_{m\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg{\}},$
@@
$\mathfrak{gl}(m|n)_{\overline{1}}=\bigg{\{}\left(\begin{array}[]{cc}0&B\\ C&0\end{array}\right)\bigg{|}B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C})\bigg{\}}.$
The parity of a homogeneous element $g\in\mathfrak{gl}(m|n)$ is denoted by $|g|$. As usual, for any $1\leq i,j\leq m+n$, let $E_{ij}$ be the $(m+n)\times(m+n)$ matrix with $1$ in the $(i,j)$-position and 00 elsewhere. Then $|E_{ij}|=|e_{i}|+|e_{j}^{\ast}|$.
@@ 2.3. Representations on $V^{\otimes d}$
On the other hand, the symmetric group $\mathfrak{S}_{d}$ acts from right on $[m+n]^{d}=\{(i_{d},\ldots,i_{1})|1\leq i_{1},\ldots,i_{d}\leq m+n\}$ as follows: for each $1\leq k<d$ and $I\in[m+n]^{d}$,
$I.s_{k}=(i_{d},\ldots,i_{k},i_{k+1},\ldots,i_{1}).$
Thus $I.\sigma=(i_{\sigma(d)},\ldots,i_{\sigma(1)})$ for $\sigma\in\mathfrak{S}_{d}$. We then have a right $\mathfrak{S}_{d}$-action on $V^{\otimes d}$ defined by
$(v_{d}\otimes\cdots\otimes v_{1}).s_{i}=(-1)^{|v_{i}|\cdot|v_{i+1}|}v_{d}\otimes\cdots\otimes v_{i}\otimes v_{i+1}\otimes\cdots\otimes v_{1}.$
@@
$(v_{3}\otimes v_{2}\otimes v_{1}).\sigma=(-1)^{|v_{2}|\cdot|v_{1}|+|v_{3}|\cdot|v_{1}|}v_{1}\otimes v_{3}\otimes v_{2}.$
@@
(2.1) $(v_{d}\otimes v_{d-1}\otimes\cdots\otimes v_{1}).\sigma=\varepsilon_{v_{d},\ldots,v_{1};\sigma}v_{\sigma(d)}\otimes v_{\sigma(d-1)}\otimes\cdots\otimes v_{\sigma(1)},$
where $\varepsilon_{v_{d},\ldots,v_{1};\sigma}$ is either $1$ or $-1$, depending on $v_{d},\ldots,v_{1}$ and $\sigma$.
@@ 2.4. Super-extensions
We define an associative product $\cdot$ on $\overline{T}$ as follows: for any $u_{k}\cdots u_{1}\otimes\tau\in\widehat{T}_{k}(V)$ and $v_{d}\cdots v_{1}\otimes\sigma\in\overline{T}_{d}(V)$,
$(u_{k}\cdots u_{1}\otimes\tau)\cdot(v_{d}\cdots v_{1}\otimes\sigma)=u_{k}\cdots u_{1}v_{d}\cdots\otimes v_{1}\otimes\tau^{\uparrow d}\sigma.$
@@
$\begin{array}[]{cccc}L(\varphi):&\overline{T}(V)&\rightarrow&\overline{T},\\[3.0pt] &\psi&\mapsto&\varphi\cdot\psi.\end{array}$
@@
$=\sum_{k=1}^{d}(-1)^{|v^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle v^{\ast},v_{k}\rangle v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes(d\ d-1\ \cdots\ k+1\ k)\sigma,$
$=\sum_{k=1}^{d}\Big{(}(v^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\big{(}(v_{d}\cdots v_{1}).(s_{k}\cdots s_{d-2}s_{d-1})\big{)}\Big{)}\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma$
@@
The operators $L(\sigma)$, $L(e_{i})$, and $L(e_{i}^{\ast})$ for $\sigma\in\mathfrak{S}_{\infty},1\leq i\leq m+n$ are elements in $\mathrm{End}_{\mathfrak{S}_{\infty}}(\overline{T}(V))$, where $\mathfrak{S}_{\infty}$ acts on $\overline{T}(V)$ by right multiplication.

### v2

@@ Abstract.
We construct a super-extension of the usual $q$-tensor algebra through super-actions of Hecke algebras. A double centralizer property is established on this extension space, which provides a modest generalization of the quantum Schur-Sergeev duality. As an application, we present a new proof of the quantum Schur-Sergeev duality.
@@ 1. Introduction
Denote by $V^{\otimes d}$ the $d$-fold of a vector space $V$. In [Sc27], Schur studied the relation between the permutation action of the symmetric group $\mathfrak{S}_{d}$ and the diagonal action of the general linear group $\mathrm{GL}(V)$ on $V^{\otimes d}$. He pointed out that the actions of these two groups on $V^{\otimes d}$ admit a double centralizer property, which is just the celebrated Schur-Weyl duality. Since its importance in representation theory, this duality has been generalized in several directions during the past decades. In the late 1980s, the Schur-Sergeev duality, which is a super version of the Schur-Weyl duality, was established by Sergeev [Ser85] (also by Berele and Regev [BR87] independently). Meanwhile, Jimbo [Jim86] introduced a duality between Hecke algebras and general linear quantum groups, which is called the Schur-Jimbo duality nowadays. In [Mi06], Mitsuhashi provided a super quantized version of the Schur-Weyl duality, which can be regarded as a common generalization of the dualities due to Sergeev and Jimbo.
@@
Usually, in order to prove these dualities, people use the double commutant theorem. For example, for the classical Schur-Weyl duality, one needs to show that $\mathbb{C}\mathrm{GL}(V)=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$. Then by the semisimplicity of the group algebra $\mathbb{C}\mathfrak{S}_{d}$ one has $\mathbb{C}\mathfrak{S}_{d}=\mathrm{End}_{\mathbb{C}\mathrm{GL}(V)}(V^{\otimes d})$. The verification of the inclusion $\mathbb{C}\mathrm{GL}(V)\subset\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$ is trivial. But the reverse direction requires further non-trivial techniques. We refer the reader to [GW03] and [CW12] for detailed proof.
@@
In [It12], Itoh generalized the usual tensor algebra $T(V)$ by considering actions of symmetric groups. He studied operators of left multiplication by vectors and linear functionals, and established a duality theorem between the algebra generated by these operators and the infinity symmetric group. As a consequence, he provided a new proof of the Schur-Weyl duality theorem. In a subsequent paper [It15], he generalized his construction to quantum case in which symmetric group actions are replaced by Hecke algebra actions. Using a similar argument, he gave a new proof of Schur-Jimbo duality.
The main purpose of this paper is to generalize Itoh’s results to the quantum super case. First, we establish the relevant construction for the super case and obtain corresponding results. Then, based on the special challenges of the quantum situation, we achieve the quantum super case through a generally similar but slightly different approach. For the $q$-tensor superspace $\mathbb{V}^{\otimes d}$, we introduce a map $\gamma$ to simplify and unify the parity calculation problems that arise in the action of the Hecke algebra $\mathscr{H}_{d}$. Next, we construct the quantum super-extension algebra
$\widetilde{T}(\mathbb{V})=\bigoplus_{d\geq 0}\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{\infty},$
and define operators $L(\phi)$, $L(v^{\ast})$ over it. Following this, we carefully verify the well-definedness of the relevant operators and compute their commutation relations. Finally, in Theorem  7.2 we demonstrate a Schur-Weyl type duality
$\mathcal{A}_{d}^{k}\curvearrowright\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}\curvearrowleft\mathscr{H}_{d+k},$
where $\mathcal{A}_{d}^{k}$ is the subalgebra of $\mathrm{End}_{\mathbb{C}(q)}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
$L(w_{d})\cdots L(w_{1})L(\sigma)L(v_{1}^{\ast})\cdots L(v_{d}^{\ast})$
with $w_{1},\ldots,w_{d}\in\mathbb{V}$, $\sigma\in\mathscr{H}_{d+k}$, and $v_{1}^{*},\ldots,v_{k}^{*}\in\mathbb{V}^{*}$. In the special case of $k=0$, we recover Mitsuhashi’s duality, and Schur-Sergeev duality is a corollary for the non-quantum case.
@@
This paper is organized as follows. In Sections 2 - 4, we construct the super-extension algebra, study the operators defined over it and prove the duality theorems for super case finally. In Sections 5 - 7, we deal with the quantum super version.
@@ Acknowledgement
We would like to thank Li Luo for many helpful discussions. Especially, he introduced the Schur-Sergeev duality and its quantum version to us which initiates this work. We would also like to thank the referee for thorough review and helpful suggestions.
@@ 2.1. Symmetric groups
For any positive integer $d$, we denote by $\mathfrak{S}_{d}$ the symmetric group of the set $d]=\{1,2,\ldots,d\}$, and by $s_{i}$, $1\leq i\leq d-1$, the transposition interchanging $i$ and $i+1$. We often use cycle notation to denote a permutation. So $\tau=(i_{1}\ i_{2}\ \cdots\ i_{k})\in\mathfrak{S}_{d}$ maps $i_{r}$ to $i_{r+1}$ for $1\leq r\leq k-1$ and $i_{k}$ to $i_{1}$, and fixes other numbers. There is a natural sequence of inclusions: $\mathfrak{S}_{1}\subset\cdots\subset\mathfrak{S}_{d}\subset\cdots$. The direct limit is denoted by $\mathfrak{S}_{\infty}$. One can view $\mathfrak{S}_{\infty}$ as the group generated by $\{s_{i}\}_{i\geq 1}$ subject to the relations:
@@
Given $k\geq 1$, we define a group homomorphism $^{\uparrow k}:\mathfrak{S}_{\infty}\rightarrow\mathfrak{S}_{\infty}$ by $s_{i}^{\uparrow k}=s_{i+k}$.
@@ 2.2. Superspaces and general linear Lie superalgebras
Let $m,n$ be two fixed positive integers, and $V=\mathbb{C}^{m|n}=\mathbb{C}^{m}\oplus\mathbb{C}^{n}$. As usual, we set $V_{\overline{0}}=\mathbb{C}^{m}$ and $V_{\overline{1}}=\mathbb{C}^{n}$. Vectors in $V_{\overline{0}}$ (resp. $V_{\overline{1}}$) are of parity 0 (resp. parity $1$). The parity of a homogeneous element $v\in V$ is denoted by $|v|$. In this section, we will denote by $e_{i}$, $1\leq i\leq m+n$, the vector in $V$ that is 1 in the $i$-th position and 0 elsewhere. Then $\{e_{1},\ldots,e_{m}\}$ and $\{e_{m+1},\ldots,e_{m+n}\}$ are bases of $V_{\overline{0}}$ and $V_{\overline{1}}$, respectively. We denote $\{e_{1}^{\ast},\ldots,e_{m+n}^{\ast}\}$ as the dual basis of $\{e_{1},\ldots,e_{m+n}\}$ in the dual space $V^{\ast}$, i.e., $e_{i}^{\ast}(e_{j})=\delta_{ij}$. Here $\delta_{ij}$ is the Kronecker symbol. Define a parity function on this dual basis by requiring that $|e_{i}^{\ast}|=|e_{i}|$ for each $1\leq i\leq m+n$. Thus the dual space $V^{\ast}=V_{\overline{0}}^{\ast}\oplus V_{\overline{1}}^{\ast}$ of $V$ is a $\mathbb{Z}_{2}$-graded vector space.
@@
$\bigg\{\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\bigg|~A\in M_{m\times m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg\}.$
@@
$\mathfrak{gl}(m|n)_{\overline{0}}=\bigg\{\left(\begin{array}[]{cc}A&0\\ 0&D\end{array}\right)\bigg|~A\in M_{m\times m}(\mathbb{C}),D\in M_{n\times n}(\mathbb{C})\bigg\},$
@@
$\mathfrak{gl}(m|n)_{\overline{1}}=\bigg\{\left(\begin{array}[]{cc}0&B\\ C&0\end{array}\right)\bigg|~B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C})\bigg\}.$
The parity of a homogeneous element $g\in\mathfrak{gl}(m|n)$ is denoted by $|g|$. As usual, for any $1\leq i,j\leq m+n$, let $E_{ij}$ be the $(m+n)\times(m+n)$ matrix with $1$ in the $(i,j)$-position and 0 elsewhere. Then $|E_{ij}|=|e_{i}|+|e_{j}^{\ast}|$.
@@ 2.3. Representations on $V^{\otimes d}$
On the other hand, the symmetric group $\mathfrak{S}_{d}$ acts from right on $m+n]^{d}=\{(i_{d},\ldots,i_{1})|1\leq i_{1},\ldots,i_{d}\leq m+n\}$ as follows: for each $1\leq k<d$ and $I=(i_{d},\ldots,i_{1})\in[m+n]^{d}$,
$I\cdot s_{k}=(i_{d},\ldots,i_{k},i_{k+1},\ldots,i_{1}).$
Thus $I\cdot\sigma=(i_{\sigma(d)},\ldots,i_{\sigma(1)})$ for $\sigma\in\mathfrak{S}_{d}$. We then have a right $\mathfrak{S}_{d}$-action on $V^{\otimes d}$ defined by
$(v_{d}\otimes\cdots\otimes v_{1})\cdot s_{i}=(-1)^{|v_{i}|\cdot|v_{i+1}|}v_{d}\otimes\cdots\otimes v_{i}\otimes v_{i+1}\otimes\cdots\otimes v_{1}.$
@@
$(v_{3}\otimes v_{2}\otimes v_{1})\cdot\sigma=(-1)^{|v_{2}|\cdot|v_{1}|+|v_{3}|\cdot|v_{1}|}v_{1}\otimes v_{3}\otimes v_{2}.$
@@
(2.1) $(v_{d}\otimes v_{d-1}\otimes\cdots\otimes v_{1})\cdot\sigma=\varepsilon_{v_{d},\ldots,v_{1};\sigma}v_{\sigma(d)}\otimes v_{\sigma(d-1)}\otimes\cdots\otimes v_{\sigma(1)},$
where $\varepsilon_{v_{d},\ldots,v_{1};\sigma}$ is either $1$ or 1-1, depending on $v_{d},\ldots,v_{1}$ and $\sigma$.
@@ 2.4. Super-extensions
We define an associative product $\cdot$ on $\overline{T}(V)$ as follows: for any $u_{k}\cdots u_{1}\otimes\tau\in\overline{T}_{k}(V)$ and $v_{d}\cdots v_{1}\otimes\sigma\in\overline{T}_{d}(V)$,
$(u_{k}\cdots u_{1}\otimes\tau)\cdot(v_{d}\cdots v_{1}\otimes\sigma)=u_{k}\cdots u_{1}v_{d}\cdots v_{1}\otimes\tau^{\uparrow d}\sigma.$
@@
$\begin{array}[]{cccc}L(\varphi):&\overline{T}(V)&\rightarrow&\overline{T}(V),\\[3.0pt] &\psi&\mapsto&\varphi\cdot\psi.\end{array}$
@@
$=\sum_{k=1}^{d}(-1)^{|v^{\ast}|(|v_{d}|+\cdots+|v_{k+1}|)}\langle v^{\ast},v_{k}\rangle v_{d}\cdots\hat{v}_{k}\cdots v_{1}\otimes(d\ d-1\ \cdots\ k+1\ k)\sigma$
$=\sum_{k=1}^{d}\Big((v^{\ast}\otimes\mathrm{id}_{V^{\otimes d-1}})\big((v_{d}\cdots v_{1})\cdot(s_{k}\cdots s_{d-2}s_{d-1})\big)\Big)\otimes s_{d-1}s_{d-2}\cdots s_{k}\sigma,$
@@
The operators $L(\sigma)$, $L(e_{i})$, and $L(e_{i}^{\ast})$ for $\sigma\in\mathfrak{S}_{\infty},1\leq i\leq m+n$ are elements in $\mathrm{End}_{\mathbb{C}\mathfrak{S}_{\infty}}(\overline{T}(V))$, where $\mathfrak{S}_{\infty}$ acts on $\overline{T}(V)$ by right multiplication.